Find the roots of the function f(x) = 3^x · (log base2(x) − 3)^5 · e^x²−3x
.
1
Expert's answer
2019-11-05T14:14:26-0500
"f(x)=3^x(\log_2(x)-3)^5e^{x^2}-3x"

Domain

"D(f): (0, \infin)"

If "0<x<8," then "\log_2(x)-3<0." Hence "f(x)<0" for "0<x<8."

Find the first derivative


"f'(x)=3^x\cdot\ln3\cdot(\log_2(x)-3)^5e^{x^2}+""+5\cdot3^x\cdot{1 \over x\ln2}(\log_2(x)-3)^4e^{x^2}+""+2x\cdot3^x(\log_2(x)-3)^5e^{x^2}-3"


Find the second derivative


"f''(x)=3^x\cdot(\ln3)^2\cdot(\log_2(x)-3)^5e^{x^2}+""+5\cdot3^x\cdot{\ln3 \over x\ln2}\cdot(\log_2(x)-3)^4e^{x^2}+""+2x\ln3\cdot3^x(\log_2(x)-3)^5e^{x^2}+""+5\cdot3^x\cdot{\ln3 \over x\ln2}(\log_2(x)-3)^4e^{x^2}-""-5\cdot3^x\cdot{1 \over x^2\ln2}(\log_2(x)-3)^4e^{x^2}+""+20\cdot3^x\cdot{\ln3 \over x^2(\ln2)^2}\cdot(\log_2(x)-3)^3e^{x^2}+""+10\cdot3^x\cdot{1 \over \ln2}(\log_2(x)-3)^4e^{x^2}+""+2\cdot3^x(\log_2(x)-3)^5e^{x^2}""+2x\cdot\ln3\cdot3^x(\log_2(x)-3)^5e^{x^2}""+10\cdot3^x\cdot{1 \over \ln2}(\log_2(x)-3)^4e^{x^2}+""+4x^2\cdot3^x(\log_2(x)-3)^5e^{x^2}"

If "x>8," then "f''(x)>0." Hence "f'(x)" increases for "x>8."

"f(8)=-24<0,"

"f'(8)=-3<0,"

"f''(8)=0."


"f(8.0000000936436)\approx-24.0000"

"f'(8.0000000936436)=0.000011>0"


"f(8.00000498410413474204)\approx-0.0000000729"

"f(8.00000498410413474205)\approx0.00000004565"


The only root


"x\approx8.00000498410413474205"